The length of a rectangle is 5 meters longer than the width. Find then length and width if the perimeter is 54 meters.
Answer by checkley71(8403)

L=W+5
2L+2W=54
2(W+5)+2W=54
2W+10+2W=54
4W=54-10
4W=44
W=44/4
W=11 ANSWER FOR THE WIDTH.
L=11+5=16 ANSWER FOR THE LENGTH.
PROOF:
2*16+2*11=54
32+22=54
54=54